# Principal Components for Neural Network Initialization

### v1

@@ Abstract
Principal Component Analysis (PCA) is a commonly used tool for dimension reduction and denoising. Therefore, it is also widely used on the data prior to training a neural network. However, this approach can complicate the explanation of explainable AI (XAI) methods for the decision of the model. In this work, we analyze the potential issues with this approach and propose Principal Components-based Initialization (PCsInit), a strategy to incorporate PCA into the first layer of a neural network via initialization of the first layer in the network with the principal components, and its two variants PCsInit-Act and PCsInit-Sub. Explanations using these strategies are as direct and straightforward as for neural networks and are simpler than using PCA prior to training a neural network on the principal components. Moreover, as will be illustrated in the experiments, such training strategies can also allow further improvement of training via backpropagation.
Index Terms:
deep learning, neural networks, classification
I Introduction
@@
In the context of neural network training, PCA offers significant advantages by decorrelating the input features, which in turn leads to a better-conditioned Hessian matrix. This is crucial because a poorly conditioned Hessian, often caused by correlated features, can result in slow convergence during gradient-based optimization, leading to inefficiencies in training [1]. By transforming the input data into a set of orthogonal principal components, PCA ensures that gradient updates during backpropagation are more stable and effective, reducing the likelihood of oscillations or excessive adjustments along certain directions [2]. As a result, the optimization process becomes more efficient, with faster convergence toward the optimal solution. This makes PCA a valuable preprocessing technique for reducing the computational cost and enhancing the performance of neural networks, particularly in high-dimensional settings where the curse of dimensionality and correlated features are common [3]. Consequently, incorporating PCA in the training pipeline can lead to significant improvements in both speed and accuracy for models trained on complex, high-dimensional datasets [4]. Therefore, it is common to train neural networks on the principal components (PCA-NN).
However, this approach also has some limitations. First, the dimensionality reduction performed by PCA leads to a loss of information, meaning the back-projected gradients in the original input space may not fully capture the network’s behavior. Second, the principal components are linear combinations of original features, which may make the resulting features harder to interpret compared to the original data. So, while one can use the PCA loading matrix to transform back to the original features requires using the PCA loading matrix, which defines the contribution of each original feature to each principal component, applying PCA prior to training a neural network as in PCA-NN complicates the explanation process due to the further transformation step in PCA. Therefore, we introduce Principal Components based Initialization (PCsInit), a novel initialization technique for the first layer of a neural network based on the principal components. In addition, we also introduce PCsInit-Act, which applies an activation layer after the principal components to increase the neural network ability of nonlinear patterns. Moreover, we also introduce PCsInit-Sub, another variant of PCsInit that compute the principal components based on the subset of the input to increase the computational efficiency for large datasets. Note that these approaches are different from Neural PCA (Principal Component Analysis), which refers to using neural networks to approximate Principal Component Analysis (PCA), a technique used for dimensionality reduction and feature extraction in machine learning.
In summary, our main contributions are as follows:
  • We introduce PCsInit, a novel method for incorporating PCA into neural networks through weight initialization rather than preprocessing, which simplifies the explainability pipeline while maintaining the benefits of PCA.
  • We propose two practical variants:
    • PCsInit-Act: Enhances nonlinear pattern recognition by adding activation functions after PCA-initialized layers
    • PCsInit-Sub: Enables efficient scaling to large datasets by computing principal components on data subsets
  • We provide a theoretical analysis showing how PCsInit improves optimization by conditioning the Hessian matrix, leading to faster convergence compared to standard initialization methods.
  • Through extensive experiments on multiple datasets, we demonstrate that PCsInit and its variants achieve comparable or better performance than traditional PCA preprocessing while offering simpler explainability
II Motivation: Explainability when PCA is applied before a neural network
@@
When PCA is applied before a neural network (PCA-NN), perturbation-based XAI methods, such as LIME [5], Occlusion [6], and Feature Permutation [7], are heavily impacted because they rely on modifying input features and observing their effect on predictions. This is because perturbing the input features instead of the principal components directly affects the SVD of the input or the covariance matrix of the input and principal components are linear combinations of multiple features. On the other hand, if perturbations are applied to principal components, the results are meaningful only in the transformed PCA space. Mapping the effects back to the original features requires using the PCA loading matrix, which defines the contribution of each original feature to each principal component. By redistributing the changes in principal components proportionally to their feature weights in the loading matrix, approximate feature-level insights can be obtained. However, this process introduces inaccuracies because the dimensionality reduction step in PCA discards some variance, and the transformation is not fully invertible. Overall, perturbation-based XAI methods are inherently less interpretable in PCA-NN pipelines because the transformations obscure the direct relationship between input features and predictions.
@@
Next, let us consider gradient-based XAI methods. In a PCA-NN pipeline, the original input data $X$ is transformed into a lower-dimensional representation $Z=W^{\top}(X-\mu)$, where $W$ is the PCA projection matrix and $\mu$ is the mean of the input data, before being passed to the neural network. Gradient-based visualization methods such as saliency maps [8] and Grad-CAM [9], which rely on backpropagating gradients from the network’s output to its input, can still be applied in this setup by mapping the gradients from the PCA-transformed space $Z$ back to the original input space $X$. This mapping can be achieved using the chain rule, where $\frac{\partial\text{output}}{\partial X}=\frac{\partial\text{output}}{\partial Z}\cdot\frac{\partial Z}{\partial X}$, and for PCA, the Jacobian $\frac{\partial Z}{\partial X}=W^{\top}$. Consequently, gradients computed in the PCA space can be back-projected into the original input space to enable visualization. However, PCA may remove or distort fine-grained input features that are crucial for effective gradient-based visualization, potentially reducing the interpretability and quality of methods like saliency maps or SmoothGrad.
@@
Next, while feature attribution methods like SHAP [10] and LIME [5] can still explain predictions in the PCA-transformed space, mapping these explanations back to the original feature space for explanation of the input requires leveraging the PCA loading matrix and comes with the limitation of approximation. To be more detailed, SHAP explains predictions by calculating feature contributions based on Shapley values. In a PCA-NN setup, SHAP can be applied directly to the PCA-transformed features $Z$. To attribute importance to the original input features $X$, the contributions to $Z$ must be back-projected to $X$ using the PCA matrix $W$. However, $W$ may not be an invertible matrix. Therefore, the explanation is not complete, but rather an approximation. A similar thing happens to LIME. In summary, these methods can attribute importance in PCA-NN systems, but additional steps are required to map contributions from the reduced-dimensional PCA space back to the original input space, and the quality of these attributions depends on how much information is preserved by PCA.
On the other hand, if we can put PCA inside the neural network via initialization based on the similarity between the multiplication in PCA and the multiplication between neural network input and weight matrix, the explanation using XAI methods will be the same as for a regular neural network without applying PCA.
III Methodologies
In this section, we will introduce motivation and the formal training process of Principal Components Analysis Initialization (PCsInit), its two variations (PCsInit-Act, PCsInit-Sub), and relevant theoretical properties.
III-A Principal Components-based Initialization (PCsInit) - Basic Ideas.
@@ III Methodologies
This section details our proposed PCsInit framework. While PCA preprocessing requires maintaining and applying the transformation matrix during inference, PCsInit incorporates this information directly into the network weights, reducing operational complexity. Additionally, PCsInit allows the network to adaptively refine these initialized weights during training, potentially capturing more nuanced feature relationships that static PCA transformation might miss. Most importantly, PCsInit simplifies the explainability pipeline by eliminating the need to back-project through a separate PCA transformation step.
The basic idea of our proposed PCsInit approach using the PCA and the neural network can be described as follows. First, let $\mathbf{X}=[x_{ij}]$, where $i=1,...,n;j=1,...,p$, be an input data matrix of $n$ observations and $p$ features. Assume that the features are centered and scaled. Then, recall that the solution of PCA can be obtained using the singular value decomposition of $\mathbf{X}$:
$\mathbf{X}=\mathbf{U}\mathbf{D}\mathbf{W}^{T}$ (1)
@@
where $\mathbf{U}$ is an $n\times p$ orthogonal matrix, $\mathbf{W}$ is a $p\times p$ orthogonal matrix, and $\mathbf{D}$ is a $p\times p$ diagonal matrix with diagonal elements $d_{1}\geq d_{2}\geq...\geq d_{p}\geq 0$. If $r$ eigenvalues are used, the projection matrix is $\mathbf{W}_{r}$, which consists of the first $r$ columns of $\mathbf{W}$. Then, the dimension-reduced version of $\mathbf{X}$ is $\mathbf{X}\mathbf{W}_{r}$. This multiplication has some similarity with the multiplication between the input $\boldsymbol{X}$ and the weight matrix $\mathbf{W}$ of the first layer in a layer of a neural network. This motivates us to bring the PCA inside the neural network by simply initializing the first layer of a neural network with the principal components of the input and using no activation function for the first layer. The width of the first layer is chosen so that n_components is retained or so that $r$ percentage of the variance explained is retained. We call this strategy Principal Components-based intialization (PCsInit).
III-B Principal Components Analysis Initialization (PCsInit)
Formally, the training process of PCsInit is described in Algorithm 1. In the first step, PCA is applied to the entire input dataset $\boldsymbol{X}$ to extract the top $r$ principal components, which are stored in the projection matrix $\mathbf{W}_{r}$. These principal components represent the directions of maximum variance in the data, allowing for an efficient lower-dimensional representation while preserving critical information.
In the second step, the matrix $\mathbf{W}_{r}$ is used as the weight matrix for the first layer of a neural network $f$. This initialization ensures that the first layer performs a transformation aligned with the most informative features of the input data, providing a strong inductive bias that can facilitate learning.
At the third step, to stabilize training and allow the deeper layers to adapt to the PCA-based initialization, the first layer is frozen, meaning its weights remain unchanged, while the remaining layers of the network are trained for $n_{\text{frozen}}$ epochs. This prevents drastic weight updates in the first layer, ensuring that the extracted principal components provide a meaningful starting point for training.
At the fourth step, after the deeper layers have sufficiently adapted, the first layer is unfrozen, allowing the entire neural network to be trained jointly for an additional $n$ epochs. This fine-tuning step enables all layers, including the first one, to update their parameters in a coordinated manner, optimizing the overall network performance and leading to better feature extraction and representation learning.
@@ III-B Principal Components Analysis Initialization (PCsInit)
The reason for the initial frozen phase and then fine-tuning is that it allows the latter layers to adapt to the PCA-based feature space, establishing stable higher-level representations. Once these layers have learned effective feature combinations, unfreezing the first layer enables fine-tuning of the initial PCA-based weights while maintaining the learned feature hierarchy. This two-phase approach prevents early disruption of the meaningful PCA structure while allowing eventual optimization of all parameters.
@@
Despite the differences, note that for the PCsInit approach, if the weights of the first layer are frozen throughout the training process then it is equivalent to applying PCA to the input data and then training a model on the principal components. Even in this case, using PCA in the PCsInit manner with the first layer frozen during training makes it easier to explain the model. Specifically, a neural network built upon principal components makes decisions based on a transformed feature space where the most significant variations in the data are captured. Since PCA reduces dimensionality by keeping only the most informative features, the network learns patterns in terms of these principal components rather than the original raw features. The decision can be explained by analyzing which principal components contributed most to the output, mapping them back to the original features, and using techniques like SHAP or sensitivity analysis. Meanwhile, PCsInit allows using SHAP and other XAI techniques directly on the input. Therefore, PCsInit offers a more straightforward explanation than explaining the decision of a neural network trained on principal components.
However, sometimes computing the principal components based on the whole input matrix can induce computational cost. Meanwhile, if the first layer is not completely frozen throughout training, it will be fine-tuned later. Therefore, it may be more computationally efficient to compute the principal components based on a subset of the input matrix instead. This is the motivation for Algorithm 2, which is a slight variation of Algorithm 1, where the principal components are computed based on a subset of the input matrix to reduce the computational expense associated with finding the principal components.
@@
While PCsInit has the first layer similar to using PCA, the activation function is not applied after the first layer. Meanwhile, it is well known that activation functions are important in neural networks because they introduce non-linearity, allowing the model to learn complex patterns and relationships in data. In addition, they can also enable efficient gradient propagation during training. This motivates us to introduce another variation of PCsInit, namely PCsInit with Activation function (PCsInit-Act), which is described in Algorithm 3. It modifies PCsInit by applying an activation function after the first layer.
In summary, PCsInit-Act introduces an activation function during the initialization process, adding an additional layer of non-linearity. This enhancement improves the model’s ability to capture complex patterns within the data. On the other hand, PCsInit-Sub focuses on computational efficiency by utilizing a subset of the input data instead of the entire dataset during initialization. As will be illustrated in the experiment section, this approach provides a lightweight yet effective alternative for handling large datasets, maintaining competitive accuracy and stability while reducing computational costs. Together, these variations improve the flexibility of the PCsInit methodology, offering improved interpretability and adaptability to diverse datasets and computational constraints.
@@
3:Add an activation function after the first layer used PCA-based initialization.
@@
III-C Theoretical analysis

### v2

@@ Abstract
Principal Component Analysis (PCA) is a commonly used tool for dimension reduction and denoising. Therefore, it is also widely used on the data prior to training a neural network. However, this approach can complicate the explanation of explainable AI (XAI) methods for the decision of the model. In this work, we analyze the potential issues with this approach and propose Principal Components-based Initialization (PCsInit), a strategy to incorporate PCA into the first layer of a neural network via initialization of the first layer in the network with the principal components, and its two variants PCsInit-Act and PCsInit-Sub. Explanations using these strategies are as direct and straightforward as for neural networks and are simpler than using PCA prior to training a neural network on the principal components. We also show that the proposed techniques posess desirable theoretical properties in terms of convergence and robustness under noisy data. Moreover, as will be illustrated in the experiments, such training strategies can also allow further improvement of training via backpropagation.
1 Introduction
@@
In the context of neural network training, PCA offers significant advantages by decorrelating the input features, which in turn leads to a better-conditioned Hessian matrix. This is crucial because a poorly conditioned Hessian, often caused by correlated features, can result in slow convergence during gradient-based optimization, leading to inefficiencies in training montavon2012neural . By transforming the input data into a set of orthogonal principal components, PCA ensures that gradient updates during backpropagation are more stable and effective, reducing the likelihood of oscillations or excessive adjustments along certain directions bishop2006pattern . As a result, the optimization process becomes more efficient, with faster convergence toward the optimal solution. This makes PCA a valuable preprocessing technique for reducing the computational cost and enhancing the performance of neural networks, particularly in high-dimensional settings where the curse of dimensionality and correlated features are common hastie2009elements . Consequently, incorporating PCA in the training pipeline can lead to significant improvements in both speed and accuracy for models trained on complex, high-dimensional datasets jolliffe2016principal . Therefore, it is common to train neural networks on the principal components (PCA-NN).
@@
However, this approach also has some limitations. When PCA is applied before a neural network (PCA-NN), perturbation-based XAI methods, such as LIME ribeiro2016should , Occlusion zeiler2014visualizing , and Feature Permutation breiman2001random , are heavily impacted because they rely on modifying input features and observing their effect on predictions. This is because perturbing the input features instead of the principal components directly affects the SVD of the input or the covariance matrix of the input and principal components are linear combinations of multiple features. Moreover, if perturbations are applied to principal components, the results are meaningful only in the transformed PCA space. Mapping the effects back to the original features requires using the PCA loading matrix, which defines the contribution of each original feature to each principal component. By redistributing the changes in principal components proportionally to their feature weights in the loading matrix, approximate feature-level insights can be obtained. However, this process introduces inaccuracies because the dimensionality reduction step in PCA discards some variance, and the transformation is not fully invertible. Overall, perturbation-based XAI methods are inherently less interpretable in PCA-NN pipelines because the transformations obscure the direct relationship between input features and predictions.
@@
Next, let us consider gradient-based XAI methods. In a PCA-NN pipeline, the original input data $X$ is transformed into a lower-dimensional representation $Z=W^{\top}(X-\mu)$, where $W$ is the PCA projection matrix and $\mu$ is the mean of the input data, before being passed to the neural network. Gradient-based visualization methods such as saliency maps kadir2001saliency and Grad-CAM selvaraju2017grad , which rely on backpropagating gradients from the network’s output to its input, can still be applied in this setup by mapping the gradients from the PCA-transformed space $Z$ back to the original input space $X$. This mapping can be achieved using the chain rule, where $\frac{\partial\text{output}}{\partial X}=\frac{\partial\text{output}}{\partial Z}\cdot\frac{\partial Z}{\partial X}$, and for PCA, the Jacobian $\frac{\partial Z}{\partial X}=W^{\top}$. Consequently, gradients computed in the PCA space can be back-projected into the original input space to enable visualization. However, PCA may remove or distort fine-grained input features that are crucial for effective gradient-based visualization, potentially reducing the interpretability and quality of methods like saliency maps or SmoothGrad.
@@
Next, while feature attribution methods like SHAP lundberg2017unified and LIME ribeiro2016should can still explain predictions in the PCA-transformed space, mapping these explanations back to the original feature space for explanation of the input requires leveraging the PCA loading matrix and comes with the limitation of approximation. Specifically, in a PCA-NN setup, SHAP can be applied directly to the PCA-transformed features $Z$. To attribute importance to the original input features $X$, the contributions to $Z$ must be back-projected to $X$ using the PCA matrix $W$. However, $W$ may not be an invertible matrix. Therefore, the explanation is not complete, but rather an approximation. A similar thing happens to LIME. In summary, these methods can attribute importance in PCA-NN systems, but additional steps are required to map contributions from the reduced-dimensional PCA space back to the original input space, and the quality of these attributions depends on how much information is preserved by PCA.
@@
These limitations motivate us to introduce Principal Components based Initialization (PCsInit), a novel initialization technique for the first layer of a neural network based on the principal components that is based on the similarity between the multiplication in PCA and the multiplication between neural network input and weight matrix. Putting PCA inside the neural network via initialization as in PCsInit allows the explanation using XAI methods will be the same as for a regular neural network without applying PCA. Also, while PCA preprocessing requires maintaining and applying the transformation matrix during inference, PCsInit incorporates this information directly into the network weights, reducing operational complexity. Additionally, PCsInit allows the network to adaptively refine these initialized weights during training, potentially capturing more nuanced feature relationships that static PCA transformation might miss. Most importantly, PCsInit simplifies the explainability pipeline by eliminating the need to back-project through a separate PCA transformation step. In addition, we also introduce PCsInit-Act, which applies an activation layer after the principal components to increase the neural network ability of nonlinear patterns. Moreover, we also introduce PCsInit-Sub, another variant of PCsInit that compute the principal components based on the subset of the input to increase the computational efficiency for large datasets. Note that these approaches are different from Neural PCA valpola2015neural , which refers to using neural networks to approximate PCA.
In summary, our contributions include (i) the introduction of PCsInit, a novel method that integrates PCA into neural networks via weight initialization; (ii) we propose two variants: PCsInit-Act, which enhances nonlinear pattern recognition by applying activation functions after the PCA-initialized layer, and PCsInit-Sub, which allows efficient scaling to large datasets by computing principal components on data subsets; (iii) a theoretical analysis demonstrating the desirable properties of the proposed approach regarding convergence and robustness under noisy data; and (iv) extensive experiments across multiple datasets confirming that PCsInit and its variants achieve performance comparable to or superior to PCA-NN while maintaining a more straightforward explainability framework.
2 Methodologies
In this section, we will introduce Principal Components Analysis Initialization (PCsInit), its two variations (PCsInit-Act and PCsInit-Sub), and relevant theoretical properties.
@@ 2 Methodologies
The main idea of PCsInit can be described as follows: First, let $\mathbf{X}=[x_{ij}]$, where $i=1,...,n;j=1,...,p$, be an input data matrix of $n$ observations and $p$ features. Assume that the features are centered and scaled. Then, recall that the solution of PCA can be obtained using the singular value decomposition of $\mathbf{X}$: $\mathbf{X}=\mathbf{U}\mathbf{D}\mathbf{W}^{T},$ where $\mathbf{U}$ is an $n\times p$ orthogonal matrix, $\mathbf{W}$ is a $p\times p$ orthogonal matrix, and $\mathbf{D}$ is a $p\times p$ diagonal matrix with diagonal elements $d_{1}\geq d_{2}\geq...\geq d_{p}\geq 0$. If $r$ eigenvalues are used, the projection matrix is $\mathbf{W}_{r}$, which consists of the first $r$ columns of $\mathbf{W}$. Then, the dimension-reduced version of $\mathbf{X}$ is $\mathbf{X}\mathbf{W}_{r}$. This multiplication has some similarity with the multiplication between the input $\boldsymbol{X}$ and the weight matrix $\mathbf{W}$ of the first layer in a layer of a neural network. This motivates us to bring the PCA inside the neural network by simply initializing the first layer of a neural network with the principal components of the input and using no activation function for the first layer. The width of the first layer is chosen so that n_components is retained or so that $r$ percentage of the variance explained is retained. More formal details are presented in section 2.1 and the theoetical properties are presented in 2.2.
2.1 Principal Components Analysis Initialization (PCsInit) and its variants
Formally, the training process of PCsInit is described in Algorithm 1. In the first step, PCA is applied to the entire input dataset $\boldsymbol{X}$ to extract the top $r$ principal components, which are stored in the projection matrix $\mathbf{W}_{r}$. These principal components represent the directions of maximum variance in the data, allowing for an efficient lower-dimensional representation while preserving critical information. In the second step, the matrix $\mathbf{W}_{r}$ is used as the weight matrix for the first layer of a neural network $f$. This initialization ensures that the first layer performs a transformation aligned with the most informative features of the input data, providing a strong inductive bias that can facilitate learning. At the third step, to stabilize training and allow the deeper layers to adapt to the PCA-based initialization, the first layer is frozen, meaning its weights remain unchanged, while the remaining layers of the network are trained for $n_{\text{frozen}}$ epochs. This prevents drastic weight updates in the first layer, ensuring that the extracted principal components provide a meaningful starting point for training. At the fourth step, after the deeper layers have sufficiently adapted, the first layer is unfrozen, allowing the entire neural network to be trained jointly for an additional $n$ epochs. This fine-tuning step enables all layers, including the first one, to update their parameters in a coordinated manner, optimizing the overall network performance and leading to better feature extraction and representation learning.
@@ 2.1 Principal Components Analysis Initialization (PCsInit) and its variants
The purpose of the initial frozen phase and then fine-tuning is to allow the later layers to adapt to the PCA-based feature space, establishing stable higher-level representations. Once these layers have learned effective feature combinations, unfreezing the first layer enables fine-tuning of the initial PCA-based weights while maintaining the learned feature hierarchy. This two-phase approach prevents early disruption of the meaningful PCA structure while allowing eventual optimization of all parameters.
@@
Despite the differences, note that for the PCsInit approach, if the weights of the first layer are frozen throughout the training process then its performance is the same as applying PCA to the input data and then training a model on the principal components. This is because the first layer can be seen as the PCA-conducting process and the subsequence layers act as the regular neural network layers. Even in this case, using PCA in the PCsInit manner with the first layer frozen during training makes it easier to explain the model. Specifically, a neural network built upon principal components makes decisions based on a transformed feature space where the most significant variations in the data are captured. Since PCA reduces dimensionality by keeping only the most informative features, the network learns patterns in terms of these principal components rather than the original raw features. The decision can be explained by analyzing which principal components contributed most to the output, mapping them back to the original features, and using techniques like SHAP or sensitivity analysis. Meanwhile, PCsInit allows using SHAP and other XAI techniques directly on the input. Therefore, PCsInit offers a more straightforward explanation than explaining the decision of a neural network trained on principal components.
However, computing the principal components based on the whole input matrix can induce computational cost. Meanwhile, if the first layer is not completely frozen throughout training, it will be fine-tuned later. Therefore, it may be more computationally efficient to compute the principal components based on a subset of the input matrix instead. This motivates us to introduce PCsInit based on a SUBset of input data (PCsInit-Sub) as described in Algorithm 2, which is a slight variation of PCsInit, where the principal components are computed based on a subset of the input matrix to reduce the computational expense associated with finding the principal components.
@@
While the first layer of PCsInit is similar to using PCA, the activation function is not applied after the first layer. Meanwhile, activation functions are known to be important in neural networks as they introduce non-linearity, allowing the model to learn complex patterns and relationships in data. In addition, they can also enable efficient gradient propagation during training. This motivates us to introduce PCsInit with Activation function (PCsInit-Act), another variation of PCsInit, which is described in Algorithm 3. It modifies PCsInit by applying an activation function after the first layer, to improve the model’s ability to capture non-linear patterns within the data. On the other hand, PCsInit-Sub focuses on computational efficiency by utilizing a subset of the input data instead of the entire dataset during initialization. As will be illustrated in the experiment section, this approach provides a lightweight yet effective alternative for handling large datasets, maintaining competitive accuracy and stability while reducing computational costs. Together, these variations improve the flexibility of the PCsInit methodology, offering improved interpretability and adaptability to diverse datasets and computational constraints.
@@
3:Add an activation function (ReLU, He,…) after the first layer used PCA-based initialization.
@@
2.2 Theoretical analysis
In this section, we provide some theoretical analysis regarding the convergence and noise propagation of the PCsInit family. Without further comments, the norm used in this paper is Euclidean norm.
2.2.1 Convergence analysis